A proton is projected with a velocity of 3 × 10^6 m/s perpendicular to a uniform magnetic field of 0.6 T. Find the acceleration of the proton.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. υ = 3 × 10^6 m/s, B = 0.6 T, m = 1.67 × 10^–27 kg F = qυB q base P = 1.6 × 10^–19 C Or, a ⃗ = F/m = qυB/m = 1.6 * 10^-19 * 3 * 10^6 * 10^-1/1.67 * 10^-27 = 17.245 * 10^13 = 1.724 * 10^4 m/s^2